the annual profit from an investment is $20000 each year fof 5 years and the cost of investment is $80000 with a salvage value of $50000.The cost of capital at this risk level is 14%. Based on the given information , the net present value of the invedtment=......
Answer #1

Net present value is calculated by discounting all the net cashflows at the rate of cost of capital which is 14% and subtracting the initial investment

Net cashflow for year 1 to year 4 = 20000 and for year 5 = 20000+50000 = 70000

So, net present value of the investmen = 20000/(1+14%)^1+20000/(1+14%)^2+20000/(1+14%)^3+20000/(1+14%)^4+70000/(1+14%)^5-80000 = 14630.0526